# GeoRecon: Graph-Level Representation Learning for 3D Molecules via Reconstruction-Based Pretraining

### v1

@@ 4.4 Ablation Study
However, when the decoder becomes deeper ($L=5$), it gains sufficient capacity to locally denoise without depending as much on the global context. As a result, increasing the noise scale may not bring further
@@
benefit, and can even lead to performance degradation due to a mismatch between input difficulty and the model’s reliance on global structure. The results suggest that shallower decoders rely more on graph-level guidance, supporting our core hypothesis about the need for global conditioning.
5 Related Work
Self-Supervised Denoising Pretraining Self-supervised denoising has emerged as a principled means to leverage unlabeled molecular conformations. Coord pretraining (Zaidi et al., 2023) demonstrates that training a GNN to remove Gaussian perturbations from equilibrium structures is mathematically equivalent to learning the underlying molecular force field, markedly improving downstream QM9 performance under limited labels. Frad (Feng et al., 2023) furthers this concept by jointly perturbing dihedral angles and Cartesian coordinates, better emulating realistic conformational variations. SliDe (Ni et al., 2024) augments denoising tasks with classical force-field priors (bond lengths, angles, torsions), grounding the learned representation in established physical interactions. More recently, MolSpectra (Wang et al., 2025) integrates multimodal energy-level spectra via contrastive reconstruction, enriching the model’s access to discrete quantum information without external labels.
Multi-Task and Multi-Fidelity Learning Multi-task learning leverages shared structure across quantum properties or fidelity levels to improve sample efficiency and generalization. Predicting multiple QM9 targets jointly promotes embedding reuse and reduces overfitting. In the multi-fidelity setting, Ren et al. (2023) combine expensive CC labels with cheaper DFT data to achieve CC-level accuracy using far fewer CC samples. These approaches help mitigate data scarcity while balancing accuracy and cost.
Graph-Level Pretraining in Molecular Representation Learning While early GNN pretraining efforts focused on node-level tasks such as masked atom prediction (Hu et al., 2019; Hou et al., 2022; Xia et al., 2023), recent work has increasingly explored graph-level objectives to capture molecule-wide structural signals. For instance, GROVER (Rong et al., 2020) introduces motif prediction as a graph-level auxiliary task, encouraging the model to identify recurring functional substructures across molecules. MGSSL (Zhang et al., 2021) further extends this idea by generating molecular graphs motif-by-motif, injecting global inductive biases into the pretraining phase. GraphCL (You et al., 2020) and GCC (Qiu et al., 2020) apply graph-level contrastive learning via augmented graph views, promoting representation consistency at the whole-graph level.
Other approaches have leveraged cross-modal or multi-scale alignment to improve global structure modeling. GraphMVP (Liu et al., 2021a), for example, aligns 2D molecular graphs with their corresponding 3D conformers, transferring geometric awareness into 2D encoders through contrastive learning. Uni-Mol2 (Ji et al., 2024) introduces a two-track transformer architecture that jointly encodes atomic, graph-level, and spatial features during pretraining. While effective, many of these methods require additional modalities, large-scale augmentation pipelines, or complex model designs. In contrast, our method employs a single graph-level reconstruction objective—conditioned on pooled global embeddings—built directly atop coordinate denoising, achieving strong downstream performance without auxiliary data or significant overhead.
6 Conclusion, Limitations and Future Work
We introduced GeoRecon, a reconstruction-based pretraining framework that captures both local interactions and global molecular structures in 3D molecular graphs. By conditioning denoising on graph-level representations derived from clean conformations, GeoRecon effectively learns geometry-aware embeddings that are well-suited for downstream tasks. Despite not achieving the best results on every target, GeoRecon significantly improves upon the base model and narrows the gap to SOTA across diverse properties. These results highlight the complementary role of graph-level reconstruction in enhancing node-level pretraining, offering a conceptually simple yet effective design that can be integrated into a wide range of molecular learning frameworks. Furthermore, GeoRecon’s physically grounded and modular structure makes it compatible with various 3D molecular backbones.
Limitations. While GeoRecon demonstrates strong performance, several limitations remain. Our evaluation is currently limited to standard equilibrium datasets (QM9 and MD17), leaving open the question of generalization to more diverse, larger-scale, or noisy molecular systems, such as biomolecules or experimental crystal structures.
Future Work. Future directions include scaling GeoRecon to these more complex domains, integrating flexible attention-based backbones, and exploring contrastive or generative formulations of graph-level supervision. Moreover, while many graph-level pretraining methods—such as GROVER and GraphMAE—have shown strong results on biochemical property datasets (e.g., ESOL, Lipo, Malaria, CEP, Davis, and KIBA), they have not been evaluated on quantum-chemical benchmarks like QM9 or MD17. Extending our evaluation to include these biochemical datasets would enable a more holistic and cross-domain comparison of graph-level molecular pretraining strategies.
@@ 6 Conclusion, Limitations and Future Work
Appendix A Proof of equivariance between denoise and force field learning
@@
Let $\bm{x}_{i(j)}$ denote the position of atom $i$ in molecule $j$, where each conformation $\bm{x}_{i}\in\mathbb{R}^{3N}$ is mean-centered to ensure $\sum_{j}\bm{x}_{i(j)}=\bm{0}$, thereby lying in a $(3N-3)$-dimensional subspace. Denote the potential energy function as $U(\bm{x})$, which maps geometric configurations to their corresponding energy levels. The negative gradient $-\nabla_{\bm{x}}U(\bm{x})$ corresponds to the force field and serves as the prediction target in force-learning tasks. The distribution of molecular conformations $\bm{x}$ is denoted as $p_{\text{phy}}(\bm{x})$, which follows the Boltzmann distribution and takes the form $A\exp\left(-U(\bm{x})/kT\right)$ where $A$ is a normalization constant, $k$ is the Boltzmann constant, and $T$ is the temperature.
From a probabilistic perspective, the distribution of the noised coordinates $\bm{x}^{\texttt{denoise}}$ can be expressed as a marginalization over the clean coordinates:
$p(\bm{x}^{\texttt{denoise}})=\int p_{\tau}(\bm{x}^{\texttt{denoise}}|\bm{x}^{\texttt{clean}})p(\bm{x}^{\texttt{clean}})\,\mathrm{d}\bm{x}^{\texttt{clean}}$
Here, the conditional distribution $p_{\tau}(\bm{x}^{\texttt{denoise}}\mid\bm{x}^{\texttt{clean}})$ is modeled as a Gaussian in the $(3N-3)$-dimensional mean-centered subspace, with isotropic variance $\tau^{2}$. According to Vincent (2011), Denoising Score Match (DSM) is equivalant to Explicit Score Matching (ESM):
$\mathcal{L}_{\mathrm{DSM}}$ $=\mathbb{E}_{p_{\tau}(\bm{x}^{\texttt{denoise}},\bm{x}^{\texttt{clean}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{denoise}})-\frac{\partial\log p_{\tau}}{\partial\bm{x}^{\texttt{denoise}}}\|^{2}$
$=\mathbb{E}_{p_{\tau}(\bm{x}^{\texttt{denoise}},\bm{x}^{\texttt{clean}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{denoise}})-\frac{1}{\tau^{2}}(\bm{x}^{\texttt{denoise}}-\bm{x}^{\texttt{clean}})\|^{2}$
$=\mathbb{E}_{p_{\mathrm{phy}}(\bm{x}^{\texttt{denoise}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{denoise}})-(-\nabla U(\bm{x}^{\texttt{denoise}}))\|^{2}+\mathrm{Const}=\mathcal{L}_{\mathrm{ESM}}+\mathrm{Const}$
@@
Appendix B Supplement Ablation study

### v2

@@ A.1 Ablation of Task Rec and Clean
Effect of Reconstruction Noise Scale $\lambda$.
In subsection 4.1 we pointed out the influence of the reconstruction noise scale $\lambda$, which directly modulates the difficulty of the pretraining task. To further validate our conclusions, we performed additional experiments in response to Reviewer CY2D’s comments. Table 5 reports both the measured Lipschitz constants of the learned representations and the downstream performance on the QM9 $H_{\text{atom}}$ task, using a 14-layer encoder pretrained on a 10k subset of QM9 with varying $\lambda$.
Table 5: Effect of reconstruction noise scale $\lambda$ on Lipschitz constant (Lip) and downstream MAE ($\downarrow$) for the $H_{\text{atom}}$ task. Pretraining is performed on a 10k subset of QM9 with a 14-layer encoder.
$\lambda$ 1.15 1.20 1.35 1.50
Lip 64.505 59.934 59.489 62.945
$H_{\text{atom}}$ 7.0172 6.5091 6.4686 6.6525
As shown, moderate values of $\lambda$ (e.g., 1.201.201.351.35) yield both reduced Lipschitz constants and improved downstream accuracy, suggesting smoother and more transferable representations. This observation aligns with our intuition: larger $\lambda$ injects stronger noise, making reconstruction more challenging and forcing the model to capture global molecular dependencies, whereas excessively large $\lambda$ degrades stability. Notably, the $\lambda$ values used in our main experiments were not exhaustively tuned, implying that additional calibration may further enhance GeoRecon’s performance.
@@
However, when the decoder becomes deeper ($L=5$), it gains sufficient capacity to locally denoise without depending as much on the global context. As a result, increasing the noise scale may not bring further benefit, and can even lead to performance degradation due to a mismatch between input difficulty and the model’s reliance on global structure. The results suggest that shallower decoders rely more on graph-level guidance, supporting our core hypothesis about the need for global conditioning.
@@
Appendix B Method of Lipshitz Constant Analyze
We measured the spectral norm of the Jacobian PCQM4Mv2 for both GeoRecon and Coord after the same pre-training period, and defined the Local Lipschitz Constant (LLC) at a conformation $x$ as follows (Miyato et al., 2018):
$L(x)=\|J_{f}(x)P\|_{2}$
Here, $f$ denotes the encoder, and $J_{f}(x)\in\mathbb{R}^{d\times 3N}$ is the Jacobian of $f$ at $x$. The orthogonal projector $P\in\mathbb{R}^{3N\times 3N}$ onto the non‑rigid subspace is applied to remove rigid‑body degrees of freedom, so that the measured local Lipschitz constant reflects sensitivity only to physically meaningful (non‑rigid) deformations and is independent of global SE(3) motions. Following Gouk et al. (2021) and Novak et al. (2018), we approximate $L(x)$ via power iteration.
Appendix C Why Smoother Representations Improve Finetuning
Setting.
Let $(\mathcal{X},\bar{d})$ denote the space of molecular conformations modulo rigid motions, where $\bar{d}$ is the Procrustes / RMSD distance
$\bar{d}(x,x^{\prime})\;:=\;\min_{R\in\mathrm{SO}(3),\,t\in\mathbb{R}^{3}}\big\|x-(Rx^{\prime}+t)\big\|_{2}.$
Let $f:(\mathcal{X},\bar{d})\!\to\!(\mathbb{R}^{m},\|\cdot\|_{2})$ be a graph-level representation map with Lipschitz constant $L_{f}$, i.e., $\|f(x)-f(x^{\prime})\|_{2}\leq L_{f}\,\bar{d}(x,x^{\prime})$. We assume a training sample $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ with diameter $\mathrm{diam}_{\bar{d}}(S)\leq D$. The downstream loss $\ell$ is assumed to be $1$-Lipschitz in its prediction argument and bounded (or rescaled to $[0,1]$ for concentration). [^1] Boundedness is only used for the empirical-to-population concentration term. All results are stated on the SE(3)-quotient space and thus respect molecular symmetries.
C.1 Linear-Probe Analysis (Frozen Encoder)
We consider a linear probe on top of the frozen $f$: $g_{w}(z)=\langle w,z\rangle+b$ with $\|w\|_{2}\leq B$; denote $\mathcal{H}_{\mathrm{lin}}=\{x\mapsto g_{w}(f(x))\}$.
Theorem 2 (Smoother representations tighten generalization for linear probes).
With probability at least $1-\delta$, every $h\in\mathcal{H}_{\mathrm{lin}}$ satisfies
$\mathcal{R}(h)\;\leq\;\widehat{\mathcal{R}}_{S}(h)\;+\;\frac{c\,B\,L_{f}\,D}{\sqrt{n}}\;+\;\sqrt{\frac{\log(1/\delta)}{2n}},$
for an absolute constant $c$.
Sketch.
Apply the contraction lemma to $\phi_{y}(u)=\ell(y,u)-\ell(y,0)$, which is $1$-Lipschitz and satisfies $\phi_{y}(0)=0$, to obtain $\mathfrak{R}_{n}(\ell\!\circ\!\mathcal{H}_{\mathrm{lin}})\leq\mathfrak{R}_{n}(\mathcal{H}_{\mathrm{lin}})$. For $\mathcal{H}_{\mathrm{lin}}$ we have
$\mathfrak{R}_{n}(\mathcal{H}_{\mathrm{lin}})\leq\frac{B}{n}\,\mathbb{E}_{\sigma}\Big\|\sum_{i=1}^{n}\sigma_{i}f(x_{i})\Big\|_{2}.$
Pick any reference $x_{0}$ and set $c=f(x_{0})$; the bias $b$ absorbs this shift. Since $\|f(x_{i})-c\|_{2}\leq L_{f}\,\bar{d}(x_{i},x_{0})\leq L_{f}D$ for all $i$,
$\mathbb{E}_{\sigma}\Big\|\sum_{i=1}^{n}\sigma_{i}\big(f(x_{i})-c\big)\Big\|_{2}\;\leq\;\Big(\sum_{i=1}^{n}\|f(x_{i})-c\|_{2}^{2}\Big)^{1/2}\;\leq\;\sqrt{n}\,L_{f}D.$
Hence $\mathfrak{R}_{n}(\mathcal{H}_{\mathrm{lin}})\leq(BL_{f}D)/\sqrt{n}$. Plugging this into the standard Rademacher generalization bound for bounded $1$-Lipschitz losses yields equation 2. ∎
Lemma 1 (Noise robustness scaling with $L_{f}$ under a linear probe).
Let the input be perturbed by additive Gaussian noise $x\mapsto x+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2}I)$ in Cartesian coordinates (for $N$ atoms hence $3N$ dimensions). Then for any $w$ with $\|w\|_{2}\leq B$,
$\mathbb{E}_{\varepsilon}\!\left[\,\big|g_{w}(f(x+\varepsilon))-g_{w}(f(x))\big|\,\right]$ $\leq B\,L_{f}\,\mathbb{E}_{\varepsilon}\|\varepsilon\|_{2}=\Theta\big(B\,L_{f}\,\sigma\sqrt{3N}\big),$
$\mathbb{E}_{\varepsilon}\!\left[\,\big(g_{w}(f(x+\varepsilon))-g_{w}(f(x))\big)^{2}\,\right]$ $\leq B^{2}L_{f}^{2}\,\mathbb{E}_{\varepsilon}\|\varepsilon\|_{2}^{2}=3N\sigma^{2}B^{2}L_{f}^{2}.$
Sketch.
Alignment only decreases distance:
$\bar{d}(x,x+\varepsilon)=\min_{R,t}\|x-(R(x+\varepsilon)+t)\|_{2}\leq\|x-(x+\varepsilon)\|_{2}=\|\varepsilon\|_{2}.$
Thus,
$|g_{w}(f(x+\varepsilon))-g_{w}(f(x))|\leq\|w\|_{2}\,\|f(x+\varepsilon)-f(x)\|_{2}\leq BL_{f}\bar{d}(x,x+\varepsilon)\leq BL_{f}\|\varepsilon\|_{2}.$
Taking (squared) expectations with respect to $\varepsilon$ gives the claims. ∎
The inequality $\bar{d}(x,x+\varepsilon)\leq\|\varepsilon\|_{2}$ is a conservative upper bound, since the optimal alignment $(R,t)$ typically yields a much smaller distance in practice. Therefore, our result should be interpreted as a worst-case guarantee: the amplification of input perturbations cannot exceed $BL_{f}\|\varepsilon\|_{2}$, while in realistic settings the effect is usually weaker.
Implication.
Equations equation 2 and Lemma 1 show that a smaller $L_{f}$ (i.e., a smoother $f$) strictly tightens linear-probe generalization and attenuates noise amplification.
C.2 End-to-End Finetuning: Rigorous and Heuristic Routes
Consider now an end-to-end predictor $h=g\circ f$ where $g:\mathbb{R}^{m}\to\mathbb{R}$ is a small readout network (e.g., the MLPs used in our experiments). Let $L_{g}$ denote the Lipschitz constant of $g$ on the image of $f$ over the data domain. The composite Lipschitz constant satisfies
$L_{h}\;\leq\;L_{g}L_{f}.$ (1)
A rigorous route under spectral control.
Proposition 1 (Generalization and robustness with spectrally-controlled readouts).
Let $g$ be a depth-$d$ MLP with $1$-Lipschitz activations (e.g., ReLU) and weight matrices $W_{1},\dots,W_{d}$. Assume spectral norm bounds $\|W_{j}\|_{\mathrm{op}}\leq s_{j}$ and define the product of spectral norms $P:=\prod_{j=1}^{d}s_{j}$ and the complexity factor $S:=\big(\sum_{j=1}^{d}\|W_{j}\|_{F}^{2}/s_{j}^{2}\big)^{1/2}$. Let $L_{g}\leq P$ and $h=g\circ f$. Then, with probability at least $1-\delta$,
$\mathcal{R}(h)\,\leq\,\widehat{\mathcal{R}}_{S}(h)\,+\,\tilde{c}\frac{PSL_{f}D}{\sqrt{n}}\,+\,\sqrt{\frac{\log(1/\delta)}{2n}},\qquad\mathbb{E}_{\varepsilon}\!\left[\,(h(x+\varepsilon)-h(x))^{2}\,\right]\,\leq\,(L_{g}L_{f})^{2}\cdot 3N\sigma^{2},$
for a universal constant $\tilde{c}$ (independent of dimensions and model parameters).
Sketch.
By equation 1, $h$ is $L_{h}$-Lipschitz with $L_{h}\leq L_{g}L_{f}$; the noise bound follows as in Lemma 1. For generalization, apply a vector (Gaussian) contraction inequality to the composed class:
$\mathfrak{R}_{n}(\{g\circ f\})\;\leq\;C\,\Big(\prod_{j=1}^{d}s_{j}\Big)\,\frac{1}{n}\,\mathbb{E}\Big\|\sum_{i=1}^{n}\gamma_{i}f(x_{i})\Big\|\;\leq\;C\,\Big(\prod_{j=1}^{d}s_{j}\Big)\,\frac{L_{f}D}{\sqrt{n}},$
and incorporate the standard spectral/Frobenius control via $S=\big(\sum_{j=1}^{d}\|W_{j}\|_{F}^{2}/s_{j}^{2}\big)^{1/2}$ to obtain the stated $\tilde{c}\,\frac{PSL_{f}D}{\sqrt{n}}$ (constants absorbed into $\tilde{c}$). ∎
Remark.
In our experiments, we did not explicitly impose spectral norm constraints via dedicated regularization. However, common practices such as weight decay and initialization schemes tend to implicitly control the spectral norms of the weights, preventing $L_{g}$ and the complexity factor $S$ from becoming excessively large. Thus Proposition 1 should be interpreted primarily as a conceptual tool, clarifying how representation smoothness ($L_{f}$) and readout complexity jointly govern generalization.
Beyond such constrained settings, we next discuss a heuristic justification based on early-stage linearization.
A heuristic route via early-stage linearization.
While the exact Rademacher complexity of general MLPs depends on architectural details, early end-to-end training is well-approximated by a linearized model around the pretrained parameters (NTK-style local linearization). Under this viewpoint, the estimation term scales as $\tilde{\mathcal{O}}(L_{g}L_{f}D/\sqrt{n})$ with $L_{g}$ the local Lipschitz constant near initialization, and the noise bound remains governed solely by the composite Lipschitz constant $(L_{g}L_{f})$.
To empirically support our heuristic argument on early-stage linearization, we conduct a sanity-check experiment measuring both the cosine similarity between the predictions of the full model and its NTK approximation, and the alignment of parameter gradients across steps. As shown in Table 8, both metrics remain high (>0.85>0.85) during the first 300 steps, confirming that finetuning dynamics are well-approximated by a linear regime in the early stage. Although the similarity gradually decreases later, this evidence substantiates our claim that smoother encoders (smaller $L_{f}$) can benefit optimization through improved linearized dynamics.
Table 8: Sanity check for linearization regime during early finetuning. We report the cosine similarity between predictions of the full model and its linearized NTK approximation, as well as the gradient alignment, measured as the cosine similarity between parameter gradients at consecutive steps. Higher values indicate stronger validity of the linearization heuristic.
Step Range Cosine Sim. (Full vs. NTK) Grad. Alignment
0–100 0.98 1.000
100–200 0.90 0.927
200–300 0.87 0.921
As expected, the similarity metrics are highest in the very first steps and gradually decrease, which aligns with the known phenomenon that NTK approximations are most accurate during the initial phase of training.
Practical relevance.
Two considerations explain why full finetuning can match or surpass linear probes: (i) Certificate existence: the hypothesis class for full finetuning strictly contains that of linear probes (by taking $g$ to be linear and freezing $f$), so Theorem 2 remains attainable as a special case. (ii) Local linearization: in early training, dynamics are well-approximated by a linearized model around the pretrained parameters, effectively optimizing a linear head on a fixed $f$ and thus retaining the same $L_{f}$ dependence; with spectral/weight decay control, $L_{g}$ (and hence $L_{h}$) stays finite. If end-to-end finetuning underperforms compared to linear probing, this often suggests limitations in the optimization procedure’s ability to leverage the expanded hypothesis class, rather than deficiencies in the representation itself.
Conclusion.
Both the linear-probe guarantee (Theorem 2) and the spectrally-controlled end-to-end bound (Proposition 1), together with our empirical ablations, indicate that pretraining strategies which produce smoother graph-level encoders (smaller $L_{f}$) can improve finetuning generalization and reduce sensitivity to coordinate perturbations. Practically, spectral or norm-based regularization on the readout complements smaller $L_{f}$, yielding tighter overall control via $L_{h}\leq L_{g}L_{f}$.
Scope of analysis.
Our theoretical results should be interpreted at three different levels of rigor. (i) Theorem 2 provides a tight and formally complete guarantee for the case of linear probing, where the dependence on $L_{f}$ can be isolated exactly. (ii) Proposition 1 extends this guarantee to nonlinear readouts under explicit spectral norm and Frobenius norm control, which mathematically preserves the same $L_{f}$-dependence but requires additional architectural constraints. (iii) Beyond these constrained settings, inspired by Jacot et al. (2018), we provide a heuristic justification via early-stage linearization, suggesting that the benefits of smoothness $L_{f}$ are likely to extend to practical end-to-end finetuning. While the fully unconstrained nonlinear case remains an open theoretical question, our experiments in main content and Table 5 consistently support the practical value of smoother representations across both linear and nonlinear downstream models.
Limitation.
Our guarantees are rigorous for linear probes; extensions to nonlinear readouts rely on either architectural constraints or heuristic arguments, and the fully unconstrained case remains open. Nonetheless, our experiments consistently indicate that smoother representations ($L_{f}$) improve performance across both linear and nonlinear settings.
@@ Conclusion.
Appendix D Proof of equivariance between denoise and force field learning
@@
Let $\bm{x}_{i(j)}$ denote the position of atom $i$ in molecule $j$, where each conformation $\bm{x}_{i}\in\mathbb{R}^{3N}$ is mean-centered to ensure $\sum_{j}\bm{x}_{i(j)}=\bm{0}$, thereby lying in a (3N3)(3N-3)-dimensional subspace. Denote the potential energy function as $U(\bm{x})$, which maps geometric configurations to their corresponding energy levels. The negative gradient $\nabla_{\bm{x}}U(\bm{x})$ corresponds to the force field and serves as the prediction target in force-learning tasks. The distribution of molecular conformations $\bm{x}$ is denoted as $p_{\text{phy}}(\bm{x})$, which follows the Boltzmann distribution and takes the form $A\exp\left(-U(\bm{x})/kT\right)$ where $A$ is a normalization constant, $k$ is the Boltzmann constant, and $T$ is the temperature.
From a probabilistic perspective, the distribution of the noised coordinates $\bm{x}^{\texttt{nsd}}$ can be expressed as a marginalization over the clean coordinates:
$p(\bm{x}^{\texttt{nsd}})=\int p_{\tau}(\bm{x}^{\texttt{nsd}}|\bm{x}^{\texttt{cln}})p(\bm{x}^{\texttt{cln}})\,\mathrm{d}\bm{x}^{\texttt{cln}}$
Here, the conditional distribution $p_{\tau}(\bm{x}^{\texttt{nsd}}\mid\bm{x}^{\texttt{cln}})$ is modeled as a Gaussian in the (3N3)(3N-3)-dimensional mean-centered subspace, with isotropic variance $\tau^{2}$. According to Vincent (2011), Denoising Score Match (DSM) is equivalant to Explicit Score Matching (ESM):
$\mathcal{L}_{\mathrm{DSM}}$ $=\mathbb{E}_{p_{\tau}(\bm{x}^{\texttt{nsd}},\bm{x}^{\texttt{cln}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{nsd}})-\frac{\partial\log p_{\tau}}{\partial\bm{x}^{\texttt{nsd}}}\|^{2}$
$=\mathbb{E}_{p_{\tau}(\bm{x}^{\texttt{nsd}},\bm{x}^{\texttt{cln}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{nsd}})-\frac{1}{\tau^{2}}(\bm{x}^{\texttt{nsd}}-\bm{x}^{\texttt{cln}})\|^{2}$
$=\mathbb{E}_{p_{\mathrm{phy}}(\bm{x}^{\texttt{nsd}})}\|\mathrm{GNN}_{\theta}(\bm{x}^{\texttt{nsd}})-(-\nabla U(\bm{x}^{\texttt{nsd}}))\|^{2}+\mathrm{Const}=\mathcal{L}_{\mathrm{ESM}}+\mathrm{Const}$
@@
This interpretation is also supported by the general connection between denoising and score matching established in Vincent (2011).
Appendix E Ablation study on task clean and rec
To rigorously demonstrate the contributions of our reconstruction task and the clean alignment task, we conducted an additional ablation study on them. We pretrained a RecOnly Model, which set the weight of Clean task weight into 0 and other hyperparameters keep unchanges.